# supReferences

### v4

@@ 2.1 The Model
with $X_{1}\sim\eta_{1}$ and where the $Z_{t}$’s and $W_{t+1}$’s are independent $\mathcal{N}_{1}(0,1)$ random variables.
@@
We assume henceforth that $y_{t}=0$ for all $t\geq 1$ and, for all $z\in\mathbb{R}$, we let $M(z,\mathrm{d}x)=\mathcal{N}_{1}(\rho z,\sigma^{2})$, $G(z)=e^{-cz^{2}/2}$ and $Q(z,\mathrm{d}x)=M(z,\mathrm{d}x)G(x)$. With this notation and assumption on $(y_{t})_{t\geq 1}$, the two sequences of distributions $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are defined by
@@
The following proposition shows that the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ admit a well-defined limit in $\mathcal{P}(\mathbb{R})$ and that, as $t\rightarrow\infty$, the impact of the initial distribution $\eta_{1}$ on $\hat{\eta}_{t}$ and on $\eta_{t}$ vanishes exponentially fast:
@@
There exist constants $(C_{\star},\sigma^{2}_{\infty})\in(0,\infty)^{2}$ and $\epsilon_{\star}\in(0,1)$ such that
$\big{\|}\eta_{t+1}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\big{\|}\leq\|\hat{\eta}_{t}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|\leq C_{\star}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
@@
We stress that, despite the simplicity of the filtering problem that we consider in this work, little is known about the finite $N$ behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}$ of $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$. Notably, using Corollary 3 in Caffarel et al., (2025) only allows to establish that, for all functions $f$ in some class of functions (containing unbounded functions) and under the assumption that $|\rho|<1$, we have $\sup_{t\geq 1}\mathbb{E}\big{[}|\eta_{t}^{N}(f)-\eta_{t}(f)|]\leq CN^{-\beta/2}$ for some unknown constants $C$ and $\beta\in(0,1]$.
2.2 Particle filters and SQMC
In the specific case where $\eta_{1}=\mathcal{N}_{1}(\mu_{1},\sigma_{1}^{2})$ for some constants $\mu_{1}\in\mathbb{R}$ and $\sigma_{1}^{2}\in(0,\infty)$ the SSM defined in (2) is a linear Gaussian SSM and the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are sequences of Gaussian distributions whose means and variances can be computed using the Kalman filter. For other choices of initial distribution $\eta_{1}$ these two sequences are intractable and therefore need to be approximated.
Algorithm 1 describes a genetic filtering algorithm that can be used to perform this task. When in Algorithm 1 the $U_{t}^{n}$’s are independent $\mathcal{U}(0,1)^{2}$ random variables we recover the bootstrap PF with multinomial resampling, which is arguably the simplest PF algorithm. It is well-known that the variance of PF estimates can be reduced by replacing multinomial resampling by more advanced resampling schemes and, for instance, Algorithm 1 reduces to the (bootstrap) PF with ordered stratified resampling when the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables while the sets ($(\{U_{t,1}^{n}\}_{n=1}^{N})_{t\geq 1}$ used in the resampling steps are independent sets of $N$ dependent $\mathcal{U}(0,1)$ random variables (see Gerber et al.,, 2019, for more details on resampling algorithms for PFs).
Sequential quasi-Monte Carlo goes one step further than PF with stratified resampling in the de-randomization process of the vanilla PF (i.e. of the PF with multinomial resampling) by also introducing, at each time $t\geq 1$, some dependence in the $\mathcal{U}(0,1)$ random numbers $\{U_{t,2}^{n}\}_{n=1}^{N}$ used to perform the mutation step. In addition, for all $t\geq 1$ the $N$ dependent $\mathcal{U}(0,1)^{2}$ random variables $\{U_{t}^{n}\}^{N}_{n=1}$ used in SQMC spread evenly over $(0,1)^{2}$ with $\mathbb{P}$-probability one, that is, with $\mathbb{P}$-probability one these sets of random variables are such that $\sup_{t\geq 1}D^{*}_{2}(\{U_{t}^{n}\}^{N}_{n=1})\leq r_{N}$ for some sequence $r_{N}\rightarrow\infty$.
Formally, we recover SQMC when, in Algorithm 1, for all $t\geq 1$ the set $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(t,2)$-sequence in based $b$ (see Section 4.1 for a definition), and these random sets are independent. In this case, with $\mathbb{P}$-probability one we have $\sup_{t\geq 1}D^{*}_{2}(\{U_{t}^{n}\}_{n=1}^{N})=\mathcal{O}(N^{-1}\log(N)^{2})$. Moreover, there exists a finite universal constant $C$ such that, for all $t\geq 1$ and any square integrable function $f:[0,1]^{2}\rightarrow\mathbb{R}$, the variance of the estimate $\frac{1}{N}\sum_{n=1}^{N}f(U_{t}^{n})$ of $\int_{[0,1]^{2}}f(u)\mathrm{d}u$ is smaller than $C$ times the variance we would have obtained if $\{U_{t}^{n}\}_{n=1}^{N}$ was a collection of independent $\mathcal{U}(0,1)^{2}$ random numbers (see Gerber,, 2015, and references therein). This latter property is important as it ensures that if for a $p=2$ the result in (1) holds for the vanilla PF then it also holds for SQMC. On the other hand, for cost per iteration of SQMC is $\mathcal{O}(N\log(N))$ against $\mathcal{O}(N)$ for PFs.
@@ 2.2 Particle filters and SQMC
6:     let $\hat{X}_{t-1}^{n}=F^{-1}_{\hat{\eta}_{t-1}^{N}}(U_{t,1}^{n})$ , $X_{t}^{n}=F^{-1}_{M(\hat{X}_{t-1}^{n},\mathrm{d}x)}(U_{t,2}^{n})$ and $W_{t}^{n}=G(X_{t}^{n})/\sum_{m=1}^{N}G(X_{t}^{m})$
7:     let $\eta_{t}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}$ and $\hat{\eta}_{t}^{N}=\sum_{n=1}^{N}W_{t}^{n}\delta_{\{X_{t}^{n}\}}$
8:end for
@@
2.3 Time uniform bounds for deterministic filtering algorithms
In the previous subsection we focussed on stochastic versions of Algorithm 1, that is on the case where $U_{t}^{n}\sim\mathcal{U}(0,1)^{2}$ for all $n\in\{1,\dots,N\}$ and all $t\geq 1$. Alternatively, one may consider determinist versions of the algorithm in which the $U_{t}^{n}$’s are non-random. We do not necessarily advocate for the use of such filtering algorithms but, as it will be clear in what follows, studying Algorithm 1 based on deterministic elements of $(0,1)^{2}$ is a key step for analysing the time evolution of $\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ and of $\|\eta_{t}^{N}-\eta_{t}\|$ for both PF and SQMC. In particular, the following lemma will play a key role to study these two algorithms:
Lemma 1.
Let $N\geq 1$ and for all $t\geq 1$ let $\big{\{}u_{t}^{n}=(u_{t,1},u_{t,2}^{n})\big{\}}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ such that, for $i=1,2$, we have $u_{t,i}\neq a2^{-k}$ for all $k\in\{1,\dots,7\}$ and all $a\in\{0,\dots,2^{7}\}$, and such that $\sup_{t\in\{1,\dots,T\}}D^{*}(\{u_{t}^{n}\}_{n=1}^{N})<2^{-10}$ for some $T\in\mathbb{N}\cup\{\infty\}$. Consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}=\{u_{t}^{n}\}_{n=1}^{N}$ for all $t\geq 1$. Then, there exists a constant $\bar{C}\in(0,\infty)$, depending only on $(c,\sigma,\rho)$ and $\eta_{1}$, such that
$\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}}$
with $\epsilon_{\star}\in(0,1)$ as in Proposition 1 and with $\delta_{N,T}=\sup_{t\in\{1,\dots,T\}}D^{*}(\{u_{t}^{n}\}_{n=1}^{N})$.
We stress that, in the lemma, we did not aim at optimizing the assumptions on ($(\{u_{t}^{n}\}_{n=1}^{N})_{\geq 1}$ but rather at obtaining a result that holds under weak assumptions on this sequence of sets. In particular, when $\{u_{t}^{n}\}_{n=1}^{N}$ is the realization of an RQMC point set the upper bound on $D^{*}(\{u_{t}^{n}\}_{n=1}^{N})$ imposed in the lemma can be drastically reduced (see Remark 5 in the Appendix).
In practice we are usually interested in estimating the expectation of some function $f$ under the filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$ and under the predictive distributions $(\eta_{t})_{t\geq 1}$. Building on the result of Lemma 1, the following lemma will be used below to address this problem using PFs and SQMC:
Lemma 2.
Consider the set up of Lemma 1 with $T\geq 2$. Then, there exist constants $(\varrho_{1},\varrho_{2})\in(0,\infty)^{2}$, depending only on $c$, $\sigma$, $\rho$ and $\eta_{1}$, such that for any function $f\in\mathcal{C}^{1}(\mathbb{R})$ verifying $\sup_{x\in\mathbb{R}}|f(x)|e^{-\varrho_{1}x^{2}}<\infty$ there is a constant $C_{f}\in(0,\infty)$, depending only on $f$, $c$, $\sigma$, $\rho$ and $\eta_{1}$, for which we have, for all $a\in(0,\infty]$ and all $\gamma\in(0,1)$, and with $\epsilon_{\star}$ and $\bar{C}$ as in Lemma 1,
$\sup_{t\in\{2,\dots,T\}}\bigg{(}\big{|}\hat{\eta}_{t}^{N}\big{(}f\mathbbm{1}_{[-a,a]}\big{)}-\hat{\eta}_{t}(f)\big{|}\vee\big{|}\eta_{t}^{N}\big{(}f\mathbbm{1}_{[-a,a]}\big{)}-\eta_{t}(f)\big{|}\Big{)}\leq\delta_{N,T}^{\gamma/2}\frac{\bar{C}F(a)}{1-\epsilon_{\star}^{1-\gamma}}+C_{f}e^{-\varrho_{2}a^{2}}$
where $F(a)=|f(a)|+|f(-a)|+\int_{-a}^{a}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $F(a)=\lim_{z\rightarrow\infty}F(z)$ when $a=\infty$.
3 Time evolution of particle filter estimates
3.1 Almost sure behaviour of particle filter estimates
We consider the use of Algorithm 1 to approximate the sequences of distributions $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ defined in Section 2.1, where the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables. As explained in Section 2.2, this framework includes the bootstrap PF with both multinomial resampling and some more advanced resampling techniques. In fact, as it will be clear from the calculations that follow, the main result of this subsection (Proposition 2) holds for any resampling mechanism.
@@ 3.1 Almost sure behaviour of particle filter estimates
Let $N\geq 1$, $\mathcal{F}^{N}_{0}=\{0,\Omega\}$ be the trivial $\sigma$-algebra and for all $t\geq 1$ let
$\mathcal{F}^{N}_{t}=\sigma\Big{(}\big{\{}(X_{s}^{n},\hat{X}_{s}^{n}),\,n=1,\dots,N,\,\,s=1,\dots,t\Big{)}.$
Next, let $[a,b]\subset\mathbb{R}$ be an arbitrary non-empty interval and note that
$\inf_{x^{\prime}:\,\rho x^{\prime}\in[a,b]}M\big{(}x^{\prime},[a,b]^{c}\big{)}\geq 1-\Phi\big{(}2(b-a)/\sigma\big{)},\quad\inf_{x^{\prime}:\,\rho x^{\prime}\not\in[a,b]}M\big{(}x^{\prime},[a,b]^{c}\big{)}\geq\frac{1}{2}.$
Therefore, letting $\varrho=\min\big{\{}1-\Phi(2(b-a)/\sigma),1/2\big{\}}>0$, it follows that
@@
Then, using the second Borel-Cantelli lemma for dependent events given in Durrett, (2019, Theorem 4.3.4, page 225), the result in (5) implies that with $\mathbb{P}$-probability one we have $\eta_{t}^{N}([a,b])=0$ infinitely often. Since $N\geq 1$ and the interval $[a,b]$ are arbitrary, we obtain in particular that
$\mathbb{P}\big{(}\hat{\eta}_{t}^{N}\big{(}[-a,a]^{c}\big{)}=1,\,\,i.o.\big{)}=\mathbb{P}\big{(}\eta_{t}^{N}\big{(}[-a,a]^{c}\big{)}=1,\,\,i.o.\big{)}=1,\quad\forall N\geq 1,\quad\forall a\in(0,\infty).$ (6)
In words, for any number of particles $N\geq 1$ and interval $[-a,a]\subset\mathbb{R}$, with $\mathbb{P}$-probability one the set $\{X_{t}^{n}\}_{n=1}^{N}$ contains no elements in $[-a,a]$ for infinitely many time instants $t\geq 1$.
By combining (6) with Proposition 1 we readily obtain the first result announced in the introductory section:

### v5

@@ 2.1 The Model
with $X_{1}\sim\eta_{1}$ and where the $Z_{t}$’s and the $W_{t+1}$’s are independent $\mathcal{N}(0,1)$ random variables.
@@
We assume henceforth that $y_{t}=0$ for all $t\geq 1$ and, for all $z\in\mathbb{R}$, we let $M(z,\mathrm{d}x)=\mathcal{N}(\rho z,\sigma^{2})$, $G(z)=e^{-cz^{2}/2}$ and $Q(z,\mathrm{d}x)=M(z,\mathrm{d}x)G(x)$. With this notation and assumption on $(y_{t})_{t\geq 1}$, the two sequences of distributions $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are defined by
@@
The following proposition shows that the sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ admit a well-defined limit in $\mathcal{P}(\mathbb{R})$ and that, as $t\rightarrow\infty$, the impact of the initial distribution $\eta_{1}$ on $\hat{\eta}_{t}$ and on $\eta_{t}$ vanishes exponentially fast.
@@
There exist constants $(C_{\star},\sigma^{2}_{\infty})\in(0,\infty)^{2}$ and $\epsilon_{\star}\in(0,1)$, where only $C_{\star}$ dependents on $\eta_{1}$, such that
$\|\hat{\eta}_{t}-\mathcal{N}(0,\sigma^{2}_{\infty})\|\leq C_{\star}\epsilon_{\star}^{t},\quad\big\|\eta_{t}-\mathcal{N}(0,\rho^{2}\sigma^{2}_{\infty}+\sigma^{2})\big\|\leq C_{\star}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
@@
We stress that, despite the simplicity of the filtering problem that we consider in this work, little is known about the finite $N$ behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}$ of the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$. Notably, using Corollary 3 in Caffarel et al., (2025) only allows to establish that, for all functions $f$ in some class of functions (containing unbounded functions) and under the assumption that $|\rho|<1$, we have $\sup_{t\geq 1}\mathbb{E}\big[|\eta_{t}^{N}(f)-\eta_{t}(f)|]\leq CN^{-\beta/2}$ for some unknown constants $C$ and $\beta\in(0,1]$.
2.2 Particle filters and sequential-quasi Monte Carlo
In the specific case where $\eta_{1}=\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ for some constants $\mu_{1}\in\mathbb{R}$ and $\sigma_{1}^{2}\in(0,\infty)$, the SSM defined in (4) is a linear Gaussian SSM and the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are sequences of Gaussian distributions whose means and variances can be computed using the Kalman filter. For other choices of initial distribution $\eta_{1}$ these two sequences are intractable and therefore need to be approximated.
Algorithm 1 describes a genetic filtering algorithm that can be used to perform this task. When in Algorithm 1 the $U_{t}^{n}$’s are independent $\mathcal{U}(0,1)^{2}$ random variables we recover the bootstrap PF with multinomial resampling, which is arguably the simplest PF algorithm. By contrast, we recover SQMC when, in Algorithm 1, for all $t\geq 1$ the set $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(t,2)$-sequence in based $b$ (see Section 1.4 for a definition), and these random sets are independent.
@@ 2.2 Particle filters and sequential-quasi Monte Carlo
6:  let $\hat{X}_{t-1}^{n}=F^{-1}_{\hat{\eta}_{t-1}^{N}}(U_{t,1}^{n})$ [resampling]
7:  let $X_{t}^{n}=F^{-1}_{M(\hat{X}_{t-1}^{n},\mathrm{d}x)}(U_{t,2}^{n})$ [mutation]
8:  let $W_{t}^{n}=G(X_{t}^{n})/\sum_{m=1}^{N}G(X_{t}^{m})$
9:  let $\eta_{t}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}$ and $\hat{\eta}_{t}^{N}=\sum_{n=1}^{N}W_{t}^{n}\delta_{\{X_{t}^{n}\}}$
10:end for
@@
3 Main results
3.1 Almost sure behaviour of particle filters
The following proposition is the first result announced in the introductory section (see also Remark 2):
Proposition 2.
Let $N\geq 1$ and consider Algorithm 1 where $\big(\{U_{t,2}^{n}\}_{n=1}^{N}\big)_{t\geq 1}$ is a sequence of independent sets of $N$ independent $\mathcal{U}(0,1)$ random variables. Then,
$\mathbb{P}\big(\hat{\eta}_{t}^{N}\big([a,b]\big)=0,\,\,i.o.\big)=\mathbb{P}\big(\eta_{t}^{N}\big([a,b]\big)=0,\,\,i.o.\big)=1,\quad\forall[a,b]\subset\mathbb{R}$ (7)
and we have
$\mathbb{P}\big(\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=\mathbb{P}\big(\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=1,\quad\forall\kappa\in(0,1).$
Remark 2.
Using (2), it follows that if we replace the discrepancy metric $\|\cdot\|_{\mathrm{D}}$ by the Kolmogorov metric $\|\cdot\|$ then the conclusion of Proposition 2 holds only for all $\kappa\in(0,1/2)$.
3.1.1 Proof of Proposition 2
Proof of Proposition 2.
@@ 3.1.1 Proof of Proposition 2
To prove the first statement of the proposition we let $\mathcal{F}^{N}_{0}=\{0,\Omega\}$ be the trivial $\sigma$-algebra,
$\mathcal{F}^{N}_{t}=\sigma\Big(\big\{(X_{s}^{n},\hat{X}_{s}^{n}),\,n=1,\dots,N,\,\,s=1,\dots,t\Big),\quad\forall t\geq 1$
and $a,b]\subset\mathbb{R}$ be arbitrary. Then, letting $\varrho=\min\big\{1-\Phi(2(b-a)/\sigma),1/2\big\}>0$, we easily check that
$\inf_{x^{\prime}:\,\rho x^{\prime}\in[a,b]}M\big(x^{\prime},[a,b]^{c}\big)\geq 1-\Phi\big(2(b-a)/\sigma\big)\geq\varrho,\quad\inf_{x^{\prime}:\,\rho x^{\prime}\not\in[a,b]}M\big(x^{\prime},[a,b]^{c}\big)\geq\frac{1}{2}\geq\varrho$
from which we deduce that
@@
By using the second Borel-Cantelli lemma for dependent events given in Durrett, (2019, Theorem 4.3.4, page 225), it follows from (8) that with $\mathbb{P}$-probability one we have $\eta_{t}^{N}([a,b])=0$ infinitely often. Since we have the implication $\eta_{t}^{N}([a,b])=0\implies\hat{\eta}_{t}^{N}([a,b])=0$, the proof of (7) is complete.
To prove the second statement of the proposition we let $\kappa\in(0,1)$ be arbitrary, and we let $a_{\kappa}\in(0,\infty)$ and $T_{\kappa}\in\mathbb{N}$ be such that $\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])\geq\kappa$ for all $t\geq T_{\kappa}$. Remark that such constants $a_{\kappa}$ and $T_{\kappa}$ exist by Proposition 1. Then, for any $t\geq T_{\kappa}$ we have the implication
$\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0\implies\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\geq|\hat{\eta}_{t}^{N}([a_{\kappa},a_{\kappa}])-\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])|\geq\kappa$
and thus it follows from (7) that $\mathbb{P}\big(\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=1$. Following similar calculations we obtain that $\mathbb{P}\big(\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=1$, completing the prove of the proposition. ∎
3.1.2 Some remarks
The result stated in (7) is the key result of Proposition 2, the second part of the proposition following directly from (7) and Proposition 1.
Inspecting the proof of Proposition 2 reveals that the negative result (7) concerning the behaviour of PFs arises for a simple and intuitive reason. On the one hand, since for all $t\geq 1$ the uniform random variables $\{U_{t,2}^{n}\}_{n=1}^{N}$ used by a PF are independent, for any interval $a,b]\subset\mathbb{R}$ there exists a constant $\varrho>0$ such that, after each mutation step of the algorithm (Line 7 in Algorithm 1), all the particles are outside the interval $[a,b]$ with probability at least $\varrho^{N}$. On the other hand, because the different mutation steps of a PF rely on independent sets of uniform random variables, with probability one this event occurs infinity often, by a suitable version of the second Borel-Cantelli lemma.
Following the discussion in the previous paragraph, and more formally following the calculations in Section 3.1.1, it should be clear that the result in (7) is not specific to the considered filtering problem but will hold in general. Notably, the problem identified in (7) is entirely caused by the mutation steps of PF algorithms. This observation implies, for instance, that (7) still holds if for all $t\geq 1$ we replace the function $G(x)=e^{-cx^{2}/2}$ by any measurable function $G_{t}:\mathbb{R}\rightarrow(0,\infty)$. We also stress that (7) is not due to the fact that we consider an unbounded state-space. To illustrate this point let $\mathsf{X}=[-l,l]$ for some constant $l\in(0,\infty)$, and let $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ be as defined in (5)-(6) where, for all $x^{\prime}\in\mathbb{R}$, $M(x^{\prime},\mathrm{d}x)$ denote the $\mathcal{N}(\rho x^{\prime},\sigma^{2})$ distribution truncated on $\mathsf{X}$ and where $\eta_{1}$ is such that $\eta_{1}(\mathsf{X})=1$. In this context, by repeating the calculations in Section 3.1.1, it is trivial to verify that if in Algorithm 1 the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables then
$\mathbb{P}\big(\hat{\eta}_{t}^{N}\big([-a,a]\big)=0,\,\,i.o.\big)=\mathbb{P}\big(\eta_{t}^{N}\big[-a,a]\big)=0,\,\,i.o.\big)=1,\quad\forall N\geq 1,\quad\forall a\in(0,l).$
Interestingly, in the literature it has been realized that if, in Algorithm 1, the function $G$ is defined by $G(x)=\mathbbm{1}_{\mathbb{R}\setminus I}(x)$ for some compact interval $I\subset\mathbb{R}$ then it may happen that, at some time $t\geq 1$, all the particles $\{X_{t}^{n}\}_{n=1}^{N}$ fall into $I$. In this case, the particle system is said to be extinct, and some estimates for the probability that this event happens before a given time instant $T\geq 2$ are provided in Del Moral, (2004, Section 7.4.1, page 231). Surprisingly, it has not been noticed that the problem whereby all the particles $\{X_{t}^{n}\}_{n=1}^{N}$ are located in an arbitrary given compact interval $I\subset\mathbb{R}$ (i) arises independently of the definition of the function $G$, and (ii) implies that PFs cannot be used to reliably estimate filtering distributions over an infinite time horizon.
3.2 Finite time horizon and probabilistic behaviour of particle filters
By Proposition 2, as time progresses a PF will inevitably reach a time instant $\tau$ at which it will fail at estimating well the filtering and/or the predictive distribution of the SSM of interest. From a practical point of view it is of key interest to study the probability that such an undesirable event happens in a finite number $T$ of time steps.
The following theorem addresses this question:
Theorem 1.
Let $N\geq 1$ and consider Algorithm 1 where $\big(\{U_{t}^{n}\}_{n=1}^{N}\big)_{t\geq 1}$ is a sequence of independent sets of $N$ independent $\mathcal{U}(0,1)^{2}$ random variables. Then, there exists a constant $\bar{C}_{1}\in(0,\infty)$ (independent of $N$) such that, for all $q\in(0,1)$ and all $T\geq 1$, with probability at least 1q1-q we have
$\max\bigg\{\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\bigg\}\leq\bar{C}_{1}\delta_{N,T,q}^{1/2}\log\big(1+\delta_{N,T,q}^{-1/2}\big)$
with $\delta_{N,T,q}=N^{-1}\big(1+\log(T/q)\big)$.
The next result is a direct consequence of Theorem 1.
Corollary 1.
Consider the set-up of Theorem 1. Then, there exists a constant $\bar{C}^{\prime}_{1}\in(0,\infty)$ such that, for any $(q,\kappa)\in(0,1)^{2}$ and $T\geq 1$, we have
$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa\Big)\geq 1-q,\quad\forall N\geq N_{T,\kappa,q}$
with $N_{T,\kappa,q}=\bar{C}^{\prime}_{1}\big\{1+\log(1+\kappa^{-1})\big\}^{2}\kappa^{-2}\big(1+\log(T/q)\big)$.
The lower bound on $N$ provided in Corollary 1 to ensure a given (probabilistic) estimation error depends on $(T,q)$ only through the quantity $\log(T/q)$. As shown in the next proposition, this dependence in the parameters $T$ and $q$ is sharp: